OpenStudy (anonymous):

The lateral area of a cube is sixty-four square kilometers. What is its side length?

OpenStudy (bahrom7893):

that is a one big cube..

OpenStudy (bahrom7893):

How many sides does a cube have?

OpenStudy (anonymous):

ummmm six?

OpenStudy (anonymous):

yeah six

OpenStudy (anonymous):

l=8?

OpenStudy (bahrom7893):

A=6*a^2 64=6a^2 64/6=a^2

OpenStudy (bahrom7893):

a=8/sqrt(6)=8sqrt(6)/6=4sqrt(6)/3

OpenStudy (anonymous):

that equals what?

OpenStudy (anonymous):

4?

OpenStudy (bahrom7893):

No that equals 4*sqrt(6)/3

OpenStudy (anonymous):

that equals 4

OpenStudy (bahrom7893):

no that equals 4*sqrt(6)/3

OpenStudy (bahrom7893):

Don't argue with a gorilla

OpenStudy (anonymous):

what is the overall answer then i came up with 4

OpenStudy (bahrom7893):

Geez man how many times do i have to tell you? Total surface area of a cube: A=6*s^2, where s is the side Total surface area is: 64 km^2, thus: 64=6s^2, where s is the side. Solve for s: 64/6=s^2 Take square root on both sides: sqrt(64/6)=sqrt(s^2) 8/sqrt(6)=s, or s=8/sqrt(6). Rationalize the denominator: s=8*sqrt(6)/6 = 4*sqrt(6)/3 (km) <-----FINAL ANSWER

OpenStudy (anonymous):

that equals what exactly

OpenStudy (bahrom7893):

Nobody knows what that equals exactly, but roughly: = 3.26598632

OpenStudy (bahrom7893):

Got it tyto?

OpenStudy (anonymous):

monkeys are buttheads

OpenStudy (bahrom7893):

so are dogs

OpenStudy (anonymous):

arggh take that back crap-slinger